OpenStudy (anonymous):

The cost for 1 1/2 pounds of candy at Mr. Bulky’s is $7. At this rate, how much is 6 pounds of candy? a. $6 c. $42 b. $28 d. $7

OpenStudy (phi):

you know you have 1 ½ pounds per $7 you can set up a proportion of pounds over dollars = pounds over dollars \[ \frac{ 1 \frac{1}{2}}{7}= \frac{6}{x} \] can you find x ?

OpenStudy (anonymous):

my answer would be B right ? cause be is 28 ?

OpenStudy (phi):

yes. that sounds right. How did you get it ?

OpenStudy (anonymous):

someone else helped me figure out were i went wrong i was dividing the 1.5 by the 7 and i should have been multiplying i think ...:?